# Scalable and robust regression models for continuous proportional data

### v1

@@ Acknowledgement
First, consider the set of $\lambda$ values such that $M(s)^{\lambda}=\{(e^{s}-1)/s)\}^{\lambda}$ is a moment generating function of some distribution $Q_{\lambda}$. This set corresponds to the non-negative integers $\{\lambda\in\mathbb{R}\backslash\{0\}:M^{\lambda}$ is the m.g.f. of some distribution $Q_{\lambda}\}=\{1,2,\dots,\}$. To see this, from Theorem 3.1 (vi) of Jørgensen, (1986), it suffices to show that the analytic continuation of $M$ to the complex plane has at least one simple zero, which is confirmed by checking $M(z)=0$ has solutions $z=\pm 2\pi i,\pm 4\pi i,\dots$, where $i$ is an imaginary unit.
@@
$\xrightarrow{\qquad\quad}$         $\xrightarrow{\qquad\quad}$
@@
We argue that there is no reason that logit link is preferred over other link functions when dealing with proportions that are not interpreted as a probability. Even if it is interpreted as a probability, logit link does not give the exactly same interpretation as the logistic regression in terms of the odds ratio, and extra care is required. For example, if $y$ is the area of dense breast tissue divided by the whole breast area on a mammogram (Peplonska et al.,, 2012), $y$ is not a probability that a patient has breast cancer. Even if $y$ is a probability of some event, the corresponding regression model (either cobin or beta) yields
$\log\left\{\frac{E(y\mid\bm{x})}{1-E(y\mid\bm{x})}\right\}=\bm{x}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}\neq E\left\{\log\left(\frac{y}{1-y}\right)\mid\bm{x}\right\}\neq\log E\left\{\left(\frac{y}{1-y}\right)\mid\bm{x}\right\}$
thus the “log odds” should be newly defined as a quantity $\log\frac{E(y\mid\bm{x})}{1-E(y\mid\bm{x})}$, which is different from the expected log odds, and also different from the log of expected odds.
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta},\bm{u})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\eta_{i})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
@@
Next, we describe the posterior inference procedure for variable dispersion micobin regression (6). Given $\bm{\lambda}$, the dispersion submodel is a negative binomial regression model $(\lambda_{i}-1)\sim\mathrm{negbin}(2,\psi_{i})$, $\mathrm{logit}(\psi_{i})=\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma}$ Therefore, under the normal prior on $\bm{\gamma}\sim N(\bm{\mu}_{\gamma},\Sigma_{\gamma})$, we can combine Pólya-Gamma (PG) data augmentation (Polson et al.,, 2013) that leads to conditionally conjugate sampling of $\bm{\gamma}$. The resulting Gibbs sampler has 5 steps, where steps 1 to 3 are equal to Algorithm 2. Denote $\bm{d}_{i}\in\mathbb{R}^{d}$ as a covariate for the dispersion model, and $D\in\mathbb{R}^{n\times d}$ as a design matrix. The new step 4 corresponds to sampling $\omega_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{PG}(1+\lambda_{i},\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma})$ for $i=1,\dots,n$. Step 5 corresponds to sampling $(\bm{\gamma}\mid\bm{\lambda},\bm{\omega})\sim N_{q}(\bm{m}_{\gamma},V_{\gamma})$, where $V_{\gamma}^{-1}=D^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\omega_{1},\dots,\omega_{n})D+\Sigma_{\gamma}^{-1}$ and $\bm{m}_{\gamma}=V_{\gamma}\{D^{\mathrm{\scriptscriptstyle T}}(1.5-0.5\lambda_{1},\dots,1.5-0.5\lambda_{n})^{\mathrm{\scriptscriptstyle T}}+\Sigma_{\gamma}^{-1}\bm{\mu}_{\gamma}\}$.
@@
S.5 Further details of spatial regression simulation study
@@
We first describe the prior specification of dispersion parameters under three different regression models. First, for micobin regression, we consider the $\psi\sim\mathrm{Unif}(2,2)$ prior, reflecting the prior belief that $E\{\operatorname*{var}(Y)\}=0.5V(\mu)$ for a given mean $\mu=E(Y)$. For the prior of $\lambda$ for cobin regression, we consider $p(\lambda)=36\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, which is derived from the marginal distribution of $\lambda$ when $(\lambda-1)\mid\psi\sim\mathrm{Negbin}(2,\psi)$ and $\psi\sim\mathrm{Beta}(2,2)$. In practice, we truncate the support of $\lambda$ at some large upper bound $L$, where we choose $L=70$ throughout the paper.
@@
The precision parameter $\phi$ in the beta regression model (the sum of two beta shape parameters) satisfies $\operatorname*{var}(Y)=\mu(1-\mu)/(1+\phi)$. While there are several different choices of $p(\phi)$ available, we choose squared uniform distribution $\phi\sim\mathrm{Unif}(0,A)^{2}$ for some $A>0$, following the suggestion of Figueroa-Zúñiga et al., (2013). To match the prior belief of $E(\psi)=1/2$ so that $E(\operatorname*{var}(Y))=0.5V(\mu)$, since $\mu(1-\mu)/3$ and $V(\mu)$ operates on a similar scale (see Figure S.5), we choose $A=8.74$ so that the prior $\phi\sim\mathrm{Unif}(0,8.74)^{2}$ satisfies $E(3/(1+\phi))\approx 1/2$. The beta regression model is fitted with Stan version 2.32.2 along with rstan version 2.32.6, with all default options using No-U-Turn sampler.
@@
$\text{negtestLL}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\log\left\{\frac{1}{M}\sum_{m=1}^{M}p(y(s_{i}^{*})\mid\bm{\theta}^{(m)})\right\}.$
@@
The benthic macroinvertebrate multivariate index (MMI) data from the 2017 National Lakes Assessment Survey (U.S. Environmental Protection Agency,, 2022) and lake watershed covariates from LakeCat data (Hill et al.,, 2018) are joined based on the unique identifier of the lake (COMID). In Table S.2, we present three lakes that exhibit zero or very low MMI values, while Table S.3 outlines the details of the 9 selected LakeCat covariates.
Table S.2: Lakes with zero or very low MMI values (multiplied by 0.01 from the original scale). COMID: lake unique identifier used in NLA 2017 and LakeCat; Site ID: lake unique ID used in NLA;
COMID Site ID MMI longitude latitude Lake name County State
22721231 NLA17_GA-10002 0 -81.893 33.300 Brierpatch lake Richmond GA
9201925 NLA17_NC-10015 0.02 -80.263 35.116 Jones pond Anson NC
22845861 NLA17_AR-10001 0.021 -92.273 34.538 Ferguson lake Saline AR
@@
We standardized all (log-transformed) covariates to have a mean of 0 and a variance of 1 prior to running MCMC, and transformed back to the original scale after sampling. Based on these standardized covariates, we used the same prior as in the simulation study. That is, $\bm{\beta}\sim N_{p}(\bm{0},100^{2}I_{p})$ for regression coefficients, a standard half-Cauchy prior on the spatial random effect standard deviation $\sigma_{u}$, beta dispersion $\phi\sim\mathrm{Unif}(0,8.74)^{2}$, cobin dispersion $p(\lambda)\propto\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, and micobin dispersion $\psi\sim\mathrm{Beta}(2,2)$. We set the upper bound of $\lambda$ to be $L=70$.

### v2

@@ Acknowledgement
[…]
@@
[…]
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta},\bm{u})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\eta_{i})$, $i=1,\dots,n$ $\triangleright$ steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
@@
Next, we describe the posterior inference procedure for variable dispersion micobin regression (7). Given $\bm{\lambda}$, the dispersion submodel is a negative binomial regression model $(\lambda_{i}-1)\sim\mathrm{negbin}(2,\psi_{i})$, $\mathrm{logit}(\psi_{i})=\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma}$ . Therefore, under the normal prior on $\bm{\gamma}\sim N(\bm{\mu}_{\gamma},\Sigma_{\gamma})$, we can combine Pólya-Gamma (PG) data augmentation (Pillow and Scott, , 2012; Polson et al., , 2013) that leads to conditionally conjugate sampling of $\bm{\gamma}$. The resulting Gibbs sampler has 5 steps, where steps 1 to 3 are equal to Algorithm 2. Denote $\bm{d}_{i}\in\mathbb{R}^{d}$ as a covariate for the dispersion model, and $D\in\mathbb{R}^{n\times d}$ as a design matrix. The new step 4 corresponds to sampling $\omega_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{PG}(1+\lambda_{i},\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma})$ for $i=1,\dots,n$. Step 5 corresponds to sampling $(\bm{\gamma}\mid\bm{\lambda},\bm{\omega})\sim N_{q}(\bm{m}_{\gamma},V_{\gamma})$, where $V_{\gamma}^{-1}=D^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\omega_{1},\dots,\omega_{n})D+\Sigma_{\gamma}^{-1}$ and $\bm{m}_{\gamma}=V_{\gamma}\{D^{\mathrm{\scriptscriptstyle T}}(1.5-0.5\lambda_{1},\dots,1.5-0.5\lambda_{n})^{\mathrm{\scriptscriptstyle T}}+\Sigma_{\gamma}^{-1}\bm{\mu}_{\gamma}\}$.
@@
S.5 Further details of simulation studies
S.5.1 Robustness of MLE to distributional misspecification and beyond
Figure S.5: Data-generating distributions used in the MLE robustness simulation study. Density plots are shown for the 1% and 99% quantiles of the covariate distribution, illustrating variation of the response distribution across the covariate range.
Figure S.5 depicts four different data-generating distributions (beta, cobin, beta rectangular, beta mixture), where the true mean $\mu_{i}=g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})=g^{-1}(x_{i1})$ with a choice of link $g$ (cobit, logit) varies from the 1% quantile of $x_{i1}$ to the 99% quantile of $x_{i1}$. In addition to bias and RMSE, Table S.2 summarizes the results in terms of length of the 95% Wald confidence interval, its empirical coverage, and negative test log-likelihood (NTLL) across 1000 replicates, where NTLL is calculated based on separate held-out data of size $n/4$. The results show that, except when the true data-generating distribution is beta, cobin regression outperforms beta regression in both empirical coverage and NTLL. These differences become more prominent as $n$ increases, since the bias in the beta regression estimates has a larger impact on inference and prediction.
Table S.2: Analysis of beta and cobin regression MLE under four different data generating distributions: beta, cobin, beta rectangular (betarec), and mixture of beta (beta3mix). Results are based on 1000 replicates under the correct mean structure but possibly misspecified response distributions. CI: 95% Wald confidence interval, cover: empirical coverage probability, NTLL: negative test log likelihood
Method, link (data&fit) Beta data Cobin data Betarec data Beta3mix data
[…]
$n$
beta regression, cobit $100$ 0.430 95.4% -0.501 0.394 94.1% -0.551 0.472 85.6% -0.363 0.384 95.6% -0.578
$400$ 0.214 94.3% -0.515 0.197 91.3% -0.559 0.237 75.2% -0.367 0.192 89.5% -0.581
$1600$ 0.108 95.0% -0.517 0.099 79.0% -0.566 0.118 32.5% -0.378 0.096 70.7% -0.586
cobin regression, cobit $001$ 0.422 94.1% -0.473 0.384 94.0% -0.563 0.484 93.0% -0.386 0.370 95.8% -0.594
$400$ 0.209 93.8% -0.481 0.191 94.6% -0.574 0.243 91.5% -0.393 0.183 96.5% -0.598
$1600$ 0.105 94.3% -0.483 0.095 95.6% -0.580 0.122 92.0% -0.402 0.092 96.3% -0.603
beta regression, logit $100$ 0.322 95.0% -0.555 0.297 94.2% -0.599 0.360 87.9% -0.415 0.291 94.3% -0.617
$400$ 0.161 94.7% -0.568 0.148 78.9% -0.606 0.180 72.8% -0.427 0.145 80.3% -0.623
$1600$ 0.080 94.0% -0.570 0.074 32.0% -0.607 0.090 28.2% -0.431 0.073 26.9% -0.628
cobin regression, logit $100$ 0.298 85.6% -0.483 0.263 96.1% -0.625 0.344 85.3% -0.397 0.254 92.0% -0.635
$400$ 0.149 86.7% -0.498 0.130 92.8% -0.635 0.174 83.9% -0.414 0.126 94.1% -0.645
$1600$ 0.074 84.7% -0.505 0.065 95.7% -0.635 0.088 82.5% -0.418 0.063 94.2% -0.648
Monte Carlo standard errors are all less than 0.002 for CI length, 1.5% for cover, and 0.008 for NTLL across all settings.
We conduct an additional simulation study to evaluate the predictive performance of cobin regression when even the mean structure is misspecified. We fitted two different logit link models (beta, cobin) to four different datasets (beta, cobin, beta rectangular, beta mixture) generated from a cobit link model. We also performed the reverse, by fitting two different cobit link models (beta, cobin) to four different datasets (beta, cobin, beta rectangular, beta mixture) generated from a logit link model.
Furthermore, we compare with the data transformation approach by fitting a robust linear model to the transformed data on the real line and then back-transforming the prediction, which also represents the case when the mean structure is misspecified. Specifically, we applied the cobit transformation $g_{\mathrm{cobit}}:(0,1)\to\mathbb{R}$ to the data generated under a logit link (across four different response distributions) and fitted a robust linear model, and also applied the logit transformation to data generated under a cobit link (also across four different response distributions). On the transformed responses $\tilde{y}_{i}$, robust linear model is fit using the GAMLSS R package (Stasinopoulos and Rigby, , 2007) assuming Student-t errors (with 3 degrees of freedom, fixed), to obtain $\hat{\bm{\beta}}_{\mathrm{trans}}$ and $\hat{\sigma}_{\mathrm{trans}}$ from the model $\tilde{y}_{i}\mid\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}t_{3}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}_{\mathrm{trans}},\sigma_{\mathrm{trans}})$ for $i=1,\dots,n$, where $t_{3}(\mu,\sigma)$ denotes a location-scale parameterization of t distribution. Predictive performance was evaluated on the original response scale. Specifically, negative test log likelihood (NTLL) was computed using the inverse transformed $t_{3}$ density supported on $(0,1)$ evaluated at the test data, and the calculation of MSPE was also based on the inverse-transformed prediction. For example, when the logit transformation is used for data transformation, MSPE was computed using $\hat{y}_{\mathrm{test}}=1/(1+\exp(-\bm{x}_{\mathrm{test}}^{\mathrm{\scriptscriptstyle T}}\hat{\bm{\beta}}_{\mathrm{trans}}))\in(0,1)$, following standard practice in the transformation-based approach.
Table S.3 summarizes the results in terms of negative test log-likelihood (NTLL) and mean squared prediction error (MSPE) on held-out data of size $n/4$ across 500 replicates. Overall, except when the true data-generating distribution is beta, cobin regression exhibits comparable predictive performance relative to beta regression, even under link misspecification. In particular, cobin regression provides better predictive accuracy than beta regression in terms of NTLL when the data are generated from the three-point beta mixture model, which represents a more complex and multimodal scenario.
From Table S.3, the transformation-based approach generally performs worse than both beta and cobin regression. In particular, predictive accuracy notably deteriorates in two settings: (1) data from a logit link with a beta response and (2) data from a logit link with a beta rectangular response. These scenarios correspond to situations where boundary-proximate responses are more common (see Figure S.5), highlighting the well-known sensitivity of the transformation-based approach to boundary-proximate observations, a limitation that persists even when a heavy-tailed error distribution is employed.
Table S.3: Predictive performance of beta regression, cobin regression, and a data transformation approach under misspecified mean settings. Results are based on 500 simulation replicates.
Beta data Cobin data Betarec data Beta3mix data
Method $n$ NTLL MSPE NTLL MSPE NTLL MSPE NTLL MSPE
Data from cobit link beta regression, logit link $100$ -0.492 0.053 -0.538 0.055 -0.360 0.093 -0.570 0.051
$400$ -0.509 0.024 -0.555 0.029 -0.360 0.046 -0.576 0.029
$1600$ -0.513 0.016 -0.557 0.021 -0.372 0.036 -0.579 0.022
cobin regression, logit link $100$ -0.462 0.064 -0.550 0.055 -0.377 0.076 -0.585 0.051
$400$ -0.471 0.031 -0.568 0.029 -0.383 0.037 -0.592 0.027
$1600$ -0.476 0.022 -0.572 0.021 -0.392 0.024 -0.596 0.020
logit transform, linear regression with $t_{3}$ error $100$ -0.451 0.087 -0.522 0.062 -0.426 0.118 -0.544 0.057
$400$ -0.465 0.040 -0.522 0.062 -0.428 0.076 -0.550 0.023
$1600$ -0.471 0.029 -0.538 0.025 -0.436 0.066 -0.551 0.015
Data from logit link beta regression, cobit $100$ -0.552 0.059 -0.599 0.054 -0.409 0.080 -0.609 0.052
$400$ -0.557 0.028 -0.608 0.026 -0.422 0.035 -0.620 0.025
$1600$ -0.561 0.020 -0.608 0.018 -0.426 0.023 -0.628 0.018
cobin regression, cobit $100$ -0.485 0.065 -0.612 0.058 -0.397 0.081 -0.618 0.056
$400$ -0.488 0.031 -0.622 0.029 -0.408 0.037 -0.632 0.028
$1600$ -0.497 0.023 -0.622 0.023 -0.409 0.024 -0.636 0.022
cobit transform, linear regression with $t_{3}$ error $100$ -0.233 0.125 -0.488 0.124 -0.186 0.216 -0.491 0.099
$400$ -0.236 0.069 -0.502 0.084 -0.191 0.152 -0.503 0.055
$1600$ -0.247 0.059 -0.502 0.078 -0.196 0.137 -0.506 0.047
MSPE are scaled by 100. Monte Carlo standard errors of NTLL are all less than 0.012 for $n=100$, 0.006 for $n=400$, and 0.003 for $n=1600$. Monte Carlo standard errors of MSPE scaled by 100 are all less than 0.007.
S.5.2 Robustness and scalability under spatial regression models
@@
We first describe the prior specification of dispersion parameters under three different regression models. First, for micobin regression, we consider the $\psi\sim\mathrm{Beta}(2,2)$ prior, reflecting the prior belief that $E\{\operatorname*{var}(Y)\}=0.5V(\mu)$ for a given mean $\mu=E(Y)$. For the prior of $\lambda$ for cobin regression, we consider $p(\lambda)=36\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, which is derived from the marginal distribution of $\lambda$ when $(\lambda-1)\mid\psi\sim\mathrm{Negbin}(2,\psi)$ and $\psi\sim\mathrm{Beta}(2,2)$. In practice, we truncate the support of $\lambda$ at some large upper bound $L$, where we choose $L=70$ throughout the paper.
@@
The precision parameter $\phi$ in the beta regression model (the sum of two beta shape parameters) satisfies $\operatorname*{var}(Y)=\mu(1-\mu)/(1+\phi)$. While there are several different choices of $p(\phi)$ available, we choose squared uniform distribution $\phi\stackrel{{\scriptstyle d}}{{=}}U^{2},U\sim\mathrm{Unif}(0,A)$ for some $A>0$, following the suggestion of Figueroa-Zúñiga et al., (2013). To match the prior belief of $E(\psi)=1/2$ so that $E(\operatorname*{var}(Y))=0.5V(\mu)$, since $\mu(1-\mu)/3$ and $V(\mu)$ operates on a similar scale (see Figure S.4), we choose $A=8.74$ so that the prior of $\phi$ satisfies $E(3/(1+\phi))\approx 1/2$. The beta regression model is fitted with Stan version 2.32.2 along with rstan version 2.32.6, with all default options using No-U-Turn sampler.
@@
$\text{negtestLL}=-\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\log\left\{\frac{1}{M}\sum_{m=1}^{M}p(y(s_{i}^{*})\mid\bm{\theta}^{(m)})\right\}.$
@@
Table S.4: Bayesian spatial regression simulation results based on 100 replicates, under the correct mean structure but misspecified distributions.
$\rho=0.1$ (moderate) $\rho=0.2$ (strong)
[…]
beta regression $(200,50)$ 0.388 89.0% 0.366 85.0%
$(400,100)$ 0.264 85.0% 0.252 80.0%
cobin regression $(200,50)$ 0.353 93.0% 0.348 95.0%
$(400,100)$ 0.246 94.0% 0.242 90.0%
micobin regression $(200,50)$ 0.352 91.0% 0.348 97.0%
$(400,100)$ 0.244 89.0% 0.240 81.0%
Monte Carlo standard errors are all less than 0.004 for 95% CI length and 4% for coverage.
Table S.4 provides additional summaries of the spatial regression simulation results in terms of 95% credible interval lengths and empirical coverage. Among the beta, cobin, and micobin models, cobin regression achieves empirical coverage closest to the nominal 0.95 level, also suggesting that it is better suited for parameter inference under potential misspecification.
S.5.3 Robustness to the presence of outliers
Figure S.6 depicts two different data-generating distributions (beta, cobin), where the true mean $\mu_{i}=g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})=g_{\mathrm{cobit}}^{-1}(x_{i1})$ varies from the 1% quantile of $x_{i1}$ to the 99% quantile of $x_{i1}$. It also describes an outlier $y^{\circ}=10^{-3}$ and the conditional distribution of the response based on the value of $\bm{x}^{\circ}=(6,6)^{\mathrm{\scriptscriptstyle T}}$, which corresponds to the conditional mean of 0.5. It also describes representative quantile residual plots (Dunn and Smyth, , 1996) for four different models fitted to the beta data with $y^{\circ}=10^{-3}$, illustrating that an outlier exhibits poor goodness of fit under the beta and cobin regression models.
For prior specification of parameters not related to the mean, we followed the same settings as in the spatial regression simulation. Specifically, we used $p(\lambda)\propto\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots,L=70$ in the cobin model; a $\mathrm{Beta}(2,2)$ prior for $\psi$ in the micobin model; $\phi\stackrel{{\scriptstyle d}}{{=}}U^{2}$ with $U\sim\mathrm{Unif}(0,8.74)$ for the beta model; and for the beta rectangular model, $\phi\stackrel{{\scriptstyle d}}{{=}}U^{2}$ with $U\sim\mathrm{Unif}(0,8.74)$ and $\alpha\sim\mathrm{Unif}(0,1)$ independently. Computations are also carried out under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
Figure S.6: (Top) Data-generating distributions used in the outlier robustness simulation study. Density plots are shown for the 1% and 99% quantiles of the covariate distribution. Outlier density corresponds to the conditional distribution of response given $\bm{x}^{\circ}=(6,6)^{\mathrm{\scriptscriptstyle T}}$, and a triangle represents $y^{\circ}=10^{-3}$. (Bottom) Quantile residual plots from four different models fitted on beta data with an outlier $y^{\circ}=10^{-3}$, the quantile residual corresponding to the outlier is marked as a triangle.
Table S.5: Outlier robustness simulation results averaged over 100 replicates, summarized in terms of posterior mean estimates and sampling efficiency.
Method Setting $\beta_{0}^{\mathrm{true}}=-6$ $\beta_{1}^{\mathrm{true}}=1$ $\beta_{2}^{\mathrm{true}}=0$ Sampling $(\bm{\beta})$
$\hat{\beta}_{0}^{\circ}$ $\hat{\beta}_{0}$ $\hat{\beta}_{1}^{\circ}$ $\hat{\beta}_{1}$ $\hat{\beta}_{2}^{\circ}$ $\hat{\beta}_{2}$ mESS time (sec)
beta regression A1 -5.899 -5.995 0.938 0.999 -0.061 -0.004 4629.6 9.5
A2 -5.850 0.908 -0.087 4650.4 9.1
cobin regression A1 -5.993 -5.993 0.982 0.999 -0.025 -0.004 1847.2 25.0
A2 -5.994 0.982 -0.025 1846.3 23.8
betarec regression A1 -5.916 -5.955 0.986 0.992 -0.004 -0.004 4727.1 22.7
A2 -5.916 0.986 -0.004 4732.8 21.8
micobin regression A1 -5.929 -5.928 0.981 0.986 -0.011 -0.004 1282.4 73.9
A2 -5.926 0.981 -0.010 1302.9 72.8
beta regression B1 -5.835 -5.934 0.915 0.977 -0.056 0.002 4656.7 9.6
B2 -5.785 0.884 -0.083 4747.7 9.3
cobin regression B1 -6.010 -6.010 0.990 1.006 -0.018 0.003 1849.2 35.4
B2 -6.010 0.989 -0.018 1849.4 33.9
betarec regression B1 -5.822 -5.866 0.958 0.965 0.001 0.001 4408.8 24.3
B2 -5.821 0.958 0.001 4346.7 23.0
micobin regression B1 -6.003 -6.003 1.002 1.005 -0.002 0.002 1139.2 89.4
B2 -6.003 1.002 -0.002 1154.0 86.9
Monte Carlo standard errors are all less than 0.019 for $\hat{\beta}^{\circ}_{0}$ and $\hat{\beta}_{0}$, 0.005 for $\hat{\beta}^{\circ}_{0}$ and $\hat{\beta}_{0}$, 0.004 for $\hat{\beta}^{\circ}_{0}$ and $\hat{\beta}_{0}$, 43.5 for mESS, and 1.9 sec for runtime.
In addition to parameter stability and empirical coverage, Table S.5 summarizes the results in terms of individual parameter estimates (with and without $y^{\circ}$) and sampling efficiency of $\bm{\beta}$. These results further support the robustness of cobin and micobin regression methods in the presence of outliers, since beta regression exhibits substantial bias. Notably, the results suggest that the poor goodness of fit of the model to the outlying observations, which is evident in both beta and cobin regression, does not necessarily lead to degraded parameter stability, as illustrated by the stable cobin regression estimates across settings. Although the gains in sampling efficiency through Kolmogorov-Gamma augmentation are less pronounced in simpler fixed-effects models, where the runtime of beta and beta rectangular regression models implemented in Stan (excluding compilation time) is relatively short, they become more substantial in hierarchical settings, as demonstrated in the previous spatial regression simulation study.
@@
S.6.1 Data processing and detailed settings
The benthic macroinvertebrate multivariate index (MMI) data from the 2017 National Lakes Assessment Survey (U.S. Environmental Protection Agency, , 2022) and lake watershed covariates from LakeCat data (Hill et al., , 2018) are joined based on the unique identifier of the lake (COMID). Table S.6 outlines the details of the 9 selected LakeCat covariates.
@@
We standardized all (log-transformed) covariates to have a mean of 0 and a variance of 1 prior to running MCMC, and transformed back to the original scale after sampling. Based on these standardized covariates, we used the same prior as in the simulation study. That is, $\bm{\beta}\sim N_{p}(\bm{0},100^{2}I_{p})$ for regression coefficients, a standard half-Cauchy prior on the spatial random effect standard deviation $\sigma_{u}$, beta dispersion $\phi\stackrel{{\scriptstyle d}}{{=}}U^{2},U\sim\mathrm{Unif}(0,8.74)$, cobin dispersion $p(\lambda)\propto\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, and micobin dispersion $\psi\sim\mathrm{Beta}(2,2)$. We set the upper bound of $\lambda$ to be $L=70$. All computations for the MMI data analysis were performed on an Apple M1 CPU.